# Graph Neural Networks for Jamming Source Localization

### v1

@@ Appendix 0.C Further Results and Analysis
Figure 5 illustrates the impact of varying parameters, such as jammer power, network density, and shadowing effects, on localization performance for CAGE, GAT, and WCL on the static experiment data. Across all conditions, CAGE consistently achieves the lowest RMSE, highlighting its robustness. Stronger shadowing ($\sigma$) degrades performance across all methods due to increased signal uncertainty [21]. An inverse relationship between jammer power and localization accuracy emerges, likely because lower power forces nodes closer to the jammer, improving localization (see Table 1). As shown in Appendix 0.F, Table 10, the 2D graphs predominantly consists of instances with a small number of nodes. This imbalance limits the model’s exposure to such graphs, leading to degraded performance on larger graphs. The lack of sufficient representation may contribute to feature explosions, a phenomenon consistent with prior findings [3].
@@
Appendix 0.D Downsampling
In real-world scenarios involving Software Defined Radio (SDR) devices, high-frequency signal sampling results in dense, high-resolution data [19], often reaching thousands of nodes per graph, as observed in our dynamic experiments, which contain approximately 6000 nodes per instance (See Appendix 0.F Table 10). We explore downsampling techniques that preserve critical signal characteristics to mitigate computational overhead while retaining the most relevant information.
We evaluate two approaches: (1) window averaging, and (2) spatial binning with noise filtering. The window averaging method segments the data into equal-sized windows based on the maximum number of nodes. Within each window, node positions and noise values are averaged to produce a single representative node, smoothing short-term fluctuations and ensuring a uniform representation. The second approach applies spatial binning, grouping nodes within fixed 1m2 grid cells to prevent oversampling in localized regions and ensure spatial coverage. After binning, noise filtering retains only nodes with the highest noise values, under the assumption that weak power observations contribute less to localization accuracy [12]. This step enhances the informativeness of the downsampled data by prioritizing regions with higher noise levels, which are more indicative of environmental and signal propagation conditions.
The ablation results in Table 8 indicate that spatial binning with noise filtering outperforms window averaging, particularly in the later segments of the trajectory, where maintaining high-information nodes is crucial. Based on these findings, we adopt the downsampling configuration with 1000 nodes using spatial binning and noise filtering for our experiments.
Table 8: Ablation study of downsampling techniques on high-resolution dynamic path data under the LDPL model. The table compares the performance of window averaging and spatial binning with noise filtering across different numbers of retained nodes ($|V|$).
$|V|$ Window averaging Spatial binning with noise filtering
200 74.5 22.1 13.2 38.7 78.4 21.6 10.5 40.5
600 63.2 21.8 14.7 34.3 63.5 16.3 9.8 31.7
800 57.1 23.9 15.7 32.2 59.3 15.3 10.4 29.7
1000 54.6 23.4 16.5 31.4 54.5 13.7 9.8 27.4
Appendix 0.E Training Details and Hyperparameter Tuning
@@
Appendix 0.F Signal Propagation and Interference Modeling
@@
Noise Floor Modeling

### v2

@@ Appendix 0.B Further Results and Analysis
Figure 5 illustrates the impact of varying parameters, such as jammer power, network density, and shadowing effects, on localization performance for CAGE, GAT, and WCL on the static experiment data. Across all conditions, CAGE consistently achieves the lowest RMSE, highlighting its robustness. Stronger shadowing ($\sigma$) degrades performance across all methods due to increased signal uncertainty [21]. An inverse relationship between jammer power and localization accuracy emerges, likely because lower power forces nodes closer to the jammer, improving localization (see Tables 8 and 1). As shown in Appendix 0.D, Table 11, the 2D graphs predominantly consists of instances with a small number of nodes. This imbalance limits the model’s exposure to such graphs, leading to degraded performance on larger graphs. The lack of sufficient representation may contribute to feature explosions, a phenomenon consistent with prior findings [3].
@@
Appendix 0.C Training Details and Hyperparameter Tuning
@@
Appendix 0.D Signal Propagation and Interference Modeling
@@ 0.D.0.1 RSSI Modeling
0.D.0.2 Noise Floor Modeling